# A Unified Framework for Inference with General Missingness Patterns and Machine Learning Imputation

### v1

@@ 2.1.2 Prediction of missing or masked values
For each missingness pattern $k=1,\ldots,K$, suppose we have access to a prediction function $h_{k}$ trained on data external to ${\mathcal{D}}_{\text{obs}}$ where $h_{k}:\prod_{\{j:\bm{R}^{(k)}_{j}=1\}}{\mathcal{Y}}_{j}\mapsto\prod_{\{j:\bm{R}^{(k)}_{j}=0\}}{\mathcal{Y}}_{j}$. A key ingredient in our proposed method will use the following operation: in addition to filling the missing values for the subjects in the stratum of the $k$-th missingness pattern, we can also mask the entries for any fully-observed subject according to the missingness pattern $k$, i.e., $\{Y_{ij}:j\in\{j^{\prime}:R_{j^{\prime}}^{(k)}=0\}\}$ for subject $i$ with ${\mathcal{C}}_{i}=\infty$ and then use $h_{k}$ to obtain their predictions to impute these “pseudo” missing entries.
In fact, to study prediction-based inference, we will show in our theoretical analyses that it is central to generalize this notion of data masking-and-imputation by missingness pattern $k$ from the fully observed subjects (${\mathcal{C}}_{i}=\infty$) to the entire population. For each missingness pattern $k=1,\ldots,K$, let the random vector $\bm{Y}^{(k)}_{\text{imp}}=\left\{G_{k}(\bm{Y}),h_{k}\left(G_{k}(\bm{Y})\right)\right\}$ be comprised of observed components $G_{k}(\bm{Y})$ and masked-and-imputed components $h_{k}\left(G_{k}(\bm{Y})\right)$, where $\bm{Y}$ follows the population full-data distribution $\mathbb{P}_{\bm{Y}}$. We assume that $\bm{Y}^{(k)}_{\text{imp}}$ is distributed according to a population-level distribution $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$. We note two key points. First, $h_{k}$ can be deterministic or stochastic as long as there exists a fixed population distribution $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$ which we will assume in theoretical analyses (Section 3). This assumption is less tenable if the prediction mappings $\{h_{k}\}$ shift over time, e.g., due to model updates that are continually pushed into online repositories to progressively impute different subsets of data. Second, the current setup does not require distinct prediction models to be trained for each of the $K$ missingness patterns. When assuming type (c) variables exist, the predictions can be accomplished by using at most $q$ univariate base prediction models whenever a missing value is encountered. In this sense, $h_{k}$ can be viewed as invoking a subset of $\sum_{j}(1-R^{(k)}_{j})$ of the $q$ base prediction models for missingness pattern $k$. For example, the model for predicting diabetes status ($Y_{i1}$) using BMI only ($Y_{ip}$) can be used for any missingness pattern $k$ with $R^{(k)}_{1}=0$ and $R^{(k)}_{p}=1$. As a result, although the number of missingness patterns may grow exponentially with the number of variables $p$, the number of base prediction models required increases only linearly.
@@ 2.2 Inferential target
In this paper, we focus on obtaining valid inference on a $d$-dimensional parameter $\bm{\theta}$} in a Z-estimation framework based on a dataset with general missingness patterns under the MAR assumption. Z-estimation framework is a widely used framework for establishing an estimator in statistical literature. Several most commonly used statistical model such as generalized linear model and generalized estimating equations belong to the Z-estimation framework. Formally, define the true population parameter $\bm{\theta}$$^{\ast}$ as the unique solution of
@@
where $\bm{\psi}$} are known vector-valued maps to a typically $d$-dimensional space, as defined in van der Vaart, (2000). For independent and identically distributed, fully observed data $\mathcal{D}$} of size $N$, a Z-estimator $\widehat{{\bm{\theta}}}_{\text{Full}}$ is obtained by solving the following empirical estimating equations:
@@
Under certain regularity conditions introduced in van der Vaart, (2000) and detailed in Section S1.2 in the Supplementary Materials, $\widehat{{\bm{\theta}}}_{\text{Full}}$ is asymptotically normal with mean $\bm{\theta}$$^{\ast}$ and variance $\dot{{\bm{\psi}}_{Full}}^{-1}\mathbb{E}[{\bm{\psi}}{\bm{\psi}}^{\top}]\dot{{\bm{\psi}}_{Full}}^{-1\top}$, where ${\bm{\psi}}={\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})$ and $\dot{{\bm{\psi}}_{Full}}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}}{\partial{\bm{\theta}}}\Bigr{|}_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]$.
@@ 2.3 Rationale of prediction-based inference
In the presence of missing values, we will propose a general approach and provide the theoretical guarantee of valid inference in Section 3. The proposed approach couples the insights from the original outcome-based prediction-powered inference (PPI) framework (Angelopoulos et al., 2023a, ; Angelopoulos et al., 2023b, ) and the classical augmented inverse probability weighted (AIPW) estimators, of which outcome-based PPI is a special case (Gronsbell et al.,, 2024). It is useful to overview the rationale behind the original outcome-based PPI approach, which is designed to incorporate additional information from observations with missing outcomes and the predicted outcomes. Technically, they can be written as a difference between a base estimate (CCA estimate) $\widehat{{\bm{\theta}}}_{\text{CC}}$ and an estimate of zero $\widehat{\mathbf{\Delta}}$: [^2] The notation of $\widehat{\mathbf{\Delta}}$ in this paper differs from the one used in Angelopoulos et al., 2023a to denote a rectifier to correct the bias introduced by ML predictions. See Example 1. The current form is more convenient for constructing more efficient estimators and their theoretical analyses.
@@
We illustrate with a simple example based on population mean estimation of a scalar outcome $Y$ in the presence of outcome missingness introduced in Angelopoulos et al., 2023a . Suppose the data are independently and identically distributed according to a population distribution comprised of two partitions of size $m$ and $M$ with marginal mean outcome $\mu=\mathbb{E}[Y]$. For the first partition, we observe the outcome $Y^{\tt lab}_{i}$ and their corresponding ML-predicted alternatives $\widehat{Y}^{\tt lab}_{i}$: ${\mathcal{D}}_{\tt lab}=\{(Y^{\tt lab}_{i},\widehat{Y}^{\tt lab}_{i})\}^{m}_{i=1}$, which is referred to as the “labeled subset”. For the second, often larger, partition of data, we do not get to observe the outcome but can obtain their ML-predicted surrogates, i.e., ${\mathcal{D}}_{\tt unlab}=\{\widehat{Y}^{\tt unlab}_{i}\}^{M}_{i=1}$, referred to as “unlabeled subset”. The full data distribution of $Y$ between the two partitions are identical, i.e., the outcome is MCAR. For both partitions, the predictions are obtained from other observed covariates and prediction models that have been externally trained to predict $Y$ using these covariates. As such, the pair (Y,Y^)($\widehat{Y}$) has the same distribution between the labeled and unlabled partitions, although $Y$ is missing for the unlabled subset.
The CCA estimator for $\mu$ is the empirical mean based on the complete subset of data: $\widehat{\theta}_{\text{CC}}=\bar{Y}^{\tt lab}$, and the corresponding estimation of zero is computed as $\widehat{\Delta}=\bar{\widehat{Y}}^{\tt lab}-\bar{\widehat{Y}}^{\tt unlab}$ because the predicted outcomes in the labeled and unlabeled partitions share the same distribution under MCAR. By rewriting $\widehat{\theta}_{\text{PP}}=\widehat{\theta}_{\text{CC}}-\widehat{\Delta}=\bar{Y}^{\tt lab}-(\bar{\widehat{Y}}^{\tt lab}-\bar{\widehat{Y}}^{\tt unlab})=\bar{\widehat{Y}}^{\tt unlab}-\underbrace{(\bar{\widehat{Y}}^{\tt lab}-\bar{Y}^{\tt lab})}_{\tt``rectifier"}$, we have
@@
where the numerator of the first term is determined by the variance of prediction; the numerator of the second term is determined by the accuracy of the outcome prediction. $\widehat{\theta}_{\text{PP}}$ will have a small variance when $M\gg m$ to let the second term dominate and the machine learning prediction is precise enough such that the variance in the second term is small. To see this, when $M\gg m$, we will have $\mathrm{Var}(\widehat{\theta}_{\text{PP}})\approx\frac{1}{m}\mathrm{Var}(Y^{\tt lab}_{i}-\widehat{Y}^{\tt lab}_{i})<\mathrm{Var}(\widehat{\theta}_{\text{CC}})=\frac{1}{m}\mathrm{Var}(Y_{i}^{\tt lab})$ when the predictions are of good quality, or more precisely when the correlation satisfies $\text{Corr}(Y,\widehat{Y})>\frac{1}{2}\sqrt{\mathrm{Var}(Y)/\mathrm{Var}(\widehat{Y})}$.
However, it is critical to understand when $\widehat{{\bm{\theta}}}_{\text{PP}}$ is more efficient than the complete-case analysis in settings other than the mean estimation problem. Gronsbell et al., (2024) analyzed the efficiency of the $\widehat{{\bm{\theta}}}_{\text{PP}}$ by deriving its influence function under the linear regression setting, and connect it to the class of AIPW estimators. Specifically, Gronsbell et al., (2024) shows that $\widehat{{\bm{\theta}}}_{\text{PP}}$ is an AIPW estimator and is not the most efficient in its class. Although the optimal estimator is difficult to obtain, a weighted version of the PPI estimator was proposed to guarantee the efficiency improvement compared to the complete-case analysis (Gronsbell et al.,, 2024). The estimator is of the form $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}_{\text{Chen}}\widehat{\mathbf{\Delta}}$, where $\widehat{\mathbf{W}}_{\text{Chen}}$ is a tuning parameter estimated from the data, originally proposed in  Chen and Chen, (2000) under a different double-sampling application context but the same setup. Furthermore, recent modifications to the $\widehat{{\bm{\theta}}}_{\text{PP}}$ all consider multiplying $\widehat{\mathbf{\Delta}}$ with a data-driven weight to seek efficiency improvement over $\widehat{{\bm{\theta}}}_{\text{CC}}$ (Angelopoulos et al., 2023b, ; Miao et al.,, 2023; Gan et al.,, 2024).
@@ 2.4 Limitations of existing prediction-based inference methods
We now discuss the limitations of current methods, which we will address. First, existing PPI methods do not account for the missing-at-random mechanism. For example, in mean estimation, suppose $Y$ is the annual income response in a questionnaire. An auxiliary variable, the level of education, affects the missing probability of $Y$, such that respondents with lower level of education tend to skip the response to the income question, leaving their $Y$ missing. In such a case, the CCA estimator $\widehat{\theta}_{\text{CC}}$ itself will become biased, as it can only reflect the income levels of the population with higher education levels. However, although being designed as an estimate of zero, $\widehat{\Delta}$ in this circumstance can only capture the mean differences of the predicted income between the two education groups, rather than the differences in true income, providing no guaranteed consistency of the result $\widehat{\theta}_{\text{PP}}$ to the average income level of the entire population. Second, existing methods only consider missing outcome or blockwise missingness scenarios, although in practice general missing data patterns commonly appear. For example, in responding to a questionnaire, respondents may randomly skip questions, leading to non-monotone missingness in the resulting dataset; in a longitudinal study, participants may skip a visit and return in the next visit (Sun and Tchetgen Tchetgen,, 2018). In such cases, existing methods can only be carried out by giving up the incomplete covariates, or truncating the data at the point where first missingness occurred.
@@ 3.1 Method
Figure 1: Proposed statistical procedure for ensuring valid prediction-based inference in datasets with general missingness patterns under the missing at random (MAR) assumption. (a) we observe a set of samples with missing values. We assign each missingness pattern with an index. (b) for the $k$-th missingness pattern, we collect the samples with ${\mathcal{C}}_{i}=k$, and impute the missing values with ML predictions. Weighted complete-case analysis are conducted over the imputed datasets to obtain $\widehat{{\bm{\gamma}}}_{2,k}$; (c) for completely observed samples, we obtain the weighted analysis estimates $\widehat{{\bm{\theta}}}_{\text{WCC}}$. For each $k$-th missingness pattern, we mask the observed values from the fully observed samples to construct each missingness pattern. We then impute the masked values with machine learning prediction and conduct weighted complete-case analysis to obtain $\widehat{{\bm{\gamma}}}_{1,k}$.
The PS-PPI estimator is defined as the difference between the weighted CCA estimator $\widehat{{\bm{\theta}}}_{\text{WCC}}$, computed from the fully observed subset, and a weighted summation of $K$ estimators of zero $\widehat{\mathbf{\Delta}}_{k}$ derived from each missingness pattern:
@@
Here we introduce these notation in turn. First, $\widehat{{\bm{\theta}}}_{\text{WCC}}$ is the weighted complete-case analysis estimator obtained from the following weighted estimating equations:
$\sum_{i=1}^{N}{\bm{\psi}}^{(i)}_{\text{WCC}}=\sum_{i=1}^{N}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}{\bm{\psi}}\left(\bm{Y}_{i};{\bm{\theta}}\right)=\bm{0}.$ (3)
@@
where $\bm{m}$} is a vector-valued function that has the same dimension as $\bm{\omega}$} and satisfies regularity conditions detailed in Appendix S1.2.
Second, $\widehat{\mathbf{\Delta}}_{k}$ is the estimate of zero from the $k$-th missingness pattern. Specifically, we define $\widehat{\mathbf{\Delta}}_{k}=\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k}$, where $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$ are the solutions of the weighted estimating equations:
@@
respectively. That is, we first compute $\widehat{{\bm{\gamma}}}_{1,k}$ by substituting the values in the missing entries with predictions generated from $h_{k}$, and solve the weighted estimating equations based on complete-case data. Second, we compute $\widehat{{\bm{\gamma}}}_{2,k}$ by imputing the actual missing entries with the predictions from $h_{k}$ using data with $k$-th missingness pattern. We will show that such construction leads to an consistent and asymptotically normal estimation of zero; as shown in Proposition 1, the population parameter of the estimating equations ${\bm{\psi}}_{1,k}$ and ${\bm{\psi}}_{2,k}$ are both ${\bm{\gamma}}^{\ast}_{k}$. To have such proposition, we need the following assumption:
@@
The proof is detailed in S1.1. The intuition behind Proposition 1 is that, after accounting for the missing mechanism, the “pseudo” imputed dataset derived from the complete-case data used in Equation (5) is distributed as $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$, which is the same distribution as the imputed dataset from data with $k$-th missingness pattern used in Equation (6).
Finally, $\{\widehat{\mathbf{W}}_{k}\}$ are a set of tuning parameters in the form of weight matrices with dimension $d\times d$. $\{\widehat{\mathbf{W}}_{k}\}$ may depend on the observed data, and in Corollary 3.2 we provide a specific construction that guarantee efficiency gains over WCCA.
@@
We provide Theorem 1 to show that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is asymptotically normal.
@@
The proof is detailed in Section S1.2 in the Supplementary Materials. The asymptotic variance-covariance matrix in the proof of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ accounts for the uncertainty of the estimated propensity score models. Following Sun and Tchetgen Tchetgen, (2018) and Robins et al., (1994), we provide a convenient robust sandwich estimators that ignores the uncertainty from the propensity score model. Such estimators results in a slightly more conservative yet valid and implementable sandwich variance for $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ when the propensity scores are estimated; the form is shown below. Our numerical experiments in Section 4 validate efficiency advantages of the PS-PPI estimator based on this strategy.
@@
From Equation (7), $\Sigma_{\text{PS-PPI}}$ is a function of $\{\mathbf{W}_{k}\}=(\mathbf{W}_{1},...,\mathbf{W}_{K})$ based on which we may seek the optimal $\{\mathbf{W}^{\ast}_{k}\}$ such that $\Sigma_{\text{PS-PPI}}(\{\mathbf{W}^{\ast}_{k}\})\preceq\Sigma_{\text{PS-PPI}}(\{\mathbf{W}_{k}\})$ for any $\{\mathbf{W}_{k}\}$, where, for two square matrices $\mathbf{A}$ and $\mathbf{B}$, $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite; $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite. Such optimal $\{\mathbf{W}_{k}\}$ has no closed-form and may be computationally expensive to solve. However, when assuming identical $\mathbf{W}_{k}$ across $k$, we have a closed-form solution of $\{\mathbf{W}^{\ast}_{k}\}$, and the resulting $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is guaranteed to be more efficient than $\widehat{{\bm{\theta}}}_{\text{WCC}}$, as we state in Corollary 3.2 and prove in Appendix S1.3.

### v2

@@ 2.1.2 Prediction of missing or masked values
For each missingness pattern $k=1,\ldots,K$, suppose we have access to a prediction function $h_{k}$ trained on data external to ${\mathcal{D}}_{\text{obs}}$ where $h_{k}:\prod_{\{j:\bm{R}^{(k)}_{j}=1\}}{\mathcal{Y}}_{j}\mapsto\prod_{\{j:\bm{R}^{(k)}_{j}=0\}}{\mathcal{Y}}_{j}$. A key ingredient in our proposed method is the following “mask-and-impute” operation: in addition to filling the missing values for the subjects in the stratum of the $k$-th missingness pattern using $h_{k}$, we can also mask the entries for any fully-observed subject according to the missingness pattern $k$ (i.e., $\{Y_{ij}:j\in\{j^{\prime}:R_{j^{\prime}}^{(k)}=0\}\}$ for subject $i$ with ${\mathcal{C}}_{i}=\infty$), and then use the same $h_{k}$ to obtain their predictions to impute these “pseudo” missing entries.
In fact, to study prediction-based inference, we will show in our theoretical analyses that it is central to generalize this notion of data masking-and-imputation by missingness pattern $k$ from the fully observed subjects (${\mathcal{C}}_{i}=\infty$) to the entire population. For each missingness pattern $k=1,\ldots,K$, let the random vector $\bm{Y}^{(k)}_{\text{imp}}=\left\{G_{k}(\bm{Y}),h_{k}\left(G_{k}(\bm{Y})\right)\right\}$ be comprised of observed components $G_{k}(\bm{Y})$ and masked-and-imputed components $h_{k}\left(G_{k}(\bm{Y})\right)$, where $\bm{Y}$ follows the population full-data distribution $\mathbb{P}_{\bm{Y}}$. We assume that $\bm{Y}^{(k)}_{\text{imp}}$ is distributed according to a population-level distribution $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$. We note two key points. First, $h_{k}$ can be deterministic or stochastic as long as there exists a fixed population distribution $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$ which we will assume in theoretical analyses (Section 3). This assumption is less tenable if the prediction mappings $\{h_{k}\}$ shift over time, e.g., due to model updates that are continually pushed into online repositories to progressively impute different subsets of data. Second, the current setup does not require distinct prediction models to be trained for each missingness pattern. In practice, the predictions can be accomplished by using at most $q$ univariate base prediction models for type (a) and (b) variables, whenever a missing value is encountered. In this sense, $h_{k}$ can be viewed as invoking a subset of $\sum_{j}(1-R^{(k)}_{j})$ of the $q$ base prediction models for missingness pattern $k$. For example, the model for predicting diabetes status using BMI only can be used for any missingness pattern $k$ with observed BMI and missing diabetes status. As a result, although the number of missingness patterns may grow exponentially with the number of variables $p$, the number of base prediction models required increases only linearly.
@@ 2.2 Inferential target
In this paper, we focus on obtaining valid statistical inference on a $d$-dimensional parameter $\bm{\theta}$} in a Z-estimation framework based on a dataset with general missingness patterns under the MAR assumption. Z-estimation is a popular framework for defining estimands and studying properties of estimators in statistical literature. Several most commonly used statistical methods, such as generalized linear model and generalized estimating equations, can be understood via the Z-estimation framework. Formally, define the true population parameter $\bm{\theta}$$^{\ast}$ as the unique solution to
@@
where $\bm{\psi}$} are known vector-valued maps to a typically $d$-dimensional space, as defined in van2000asymptotic and ${\bm{\theta}}\in\mathbb{R}^{d}$ is the parameter attached to the population distribution $\mathbb{P}_{\bm{Y}}$. For independent and identically distributed, fully observed data $\mathcal{D}$} of size $N$, a Z-estimator $\widehat{{\bm{\theta}}}_{\textrm{Full}}$ is obtained by solving the following empirical estimating equations:
@@
Under certain regularity conditions introduced in van2000asymptotic and detailed in Section S1.2 in the Supplementary Materials, $\widehat{{\bm{\theta}}}_{\text{Full}}$ is asymptotically normal with mean $\bm{\theta}$$^{\ast}$ and variance $\dot{{\bm{\psi}}}^{-1}_{\textrm{Full}}\mathbb{E}[{\bm{\psi}}^{*}{\bm{\psi}}^{*\top}]\left(\dot{{\bm{\psi}}}^{-1}_{\textrm{Full}}\right)^{\top}$, where $\dot{{\bm{\psi}}}_{\text{Full}}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}}{\partial{\bm{\theta}}}\Bigr|_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right]$ is an invertible $d\times d$ matrix and ${\bm{\psi}}^{*}={\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})$ with $\mathbb{E}[{\bm{\psi}}^{*}{\bm{\psi}}^{*\top}]<\infty$.
@@ 2.3 Rationale of prediction-based inference
In the presence of missing values, we will propose a general approach and provide the theoretical guarantee of valid inference in Section 3. The proposed approach couples the insights from the original outcome-based prediction-powered inference (PPI) framework (angelopoulos2023prediction; angelopoulos2023PPI++) and the classical augmented inverse probability weighted (AIPW) estimators, of which outcome-based PPI is a special case (gronsbell2024another). It is useful to review the rationale behind the original outcome-based PPI approach, which is designed to incorporate additional information from observations with missing outcomes and the predicted outcomes. Technically, they can be written as a difference between a base estimate (CCA estimate) $\widehat{{\bm{\theta}}}_{\text{CC}}$ and an estimate of zero $\widehat{\mathbf{\Delta}}$: [^2] The notation of $\widehat{\mathbf{\Delta}}$ in this paper differs from the one used in angelopoulos2023prediction which denotes a rectifier to correct the bias introduced by ML predictions. The current form, in our view, is more helpful for motivating the construction of more efficient estimators and for theoretical analyses.
@@
We illustrate with a simple example based on population mean estimation of a scalar outcome $Y$ in the presence of outcome missingness introduced in angelopoulos2023prediction. Suppose the data are independently and identically distributed according to a population distribution comprised of two partitions of size $m$ and $M$ with marginal mean outcome $\mu=\mathbb{E}[Y]$. For the first partition, we observe the outcome $Y^{\tt lab}_{i}$ and their corresponding ML-predicted alternatives $\widehat{Y}^{\tt lab}_{i}$: ${\mathcal{D}}_{\tt lab}=\{(Y^{\tt lab}_{i},\widehat{Y}^{\tt lab}_{i})\}^{m}_{i=1}$, which is referred to as the “labeled subset”. For the second, often larger, partition of data, we do not get to observe the outcome but can obtain their ML-predicted surrogates, i.e., ${\mathcal{D}}_{\tt unlab}=\{\widehat{Y}^{\tt unlab}_{i}\}^{M}_{i=1}$, referred to as the “unlabeled subset”. We assume the outcome is MCAR, i.e., the full data distribution of $Y$ between the two partitions is identical. For both partitions, the predictions are obtained from other observed covariates and prediction models that have been externally trained to predict $Y$ using these covariates. As such, the pair (Y,Y^)($\widehat{Y}$) has the same distribution between the labeled and unlabeled partitions, although $Y$ is missing for the unlabeled subset.
In PPI, the CCA estimator for $\mu$ is the empirical mean based on the complete subset of data: $\widehat{\theta}_{\text{CC}}=\bar{Y}^{\tt lab}$, and the corresponding estimation of zero is computed as $\widehat{\Delta}=\bar{\widehat{Y}}^{\tt lab}-\bar{\widehat{Y}}^{\tt unlab}$ because the predicted outcomes in the labeled and unlabeled partitions share the same distribution under MCAR. By rewriting $\widehat{\theta}_{\text{PP}}=\widehat{\theta}_{\text{CC}}-\widehat{\Delta}=\bar{Y}^{\tt lab}-(\bar{\widehat{Y}}^{\tt lab}-\bar{\widehat{Y}}^{\tt unlab})=\bar{\widehat{Y}}^{\tt unlab}-\underbrace{(\bar{\widehat{Y}}^{\tt lab}-\bar{Y}^{\tt lab})}_{\tt``rectifier"}$, we have
@@
where the denominator of the first term is determined by the sample size of the unlabeled subset $M$, while the numerator of the second term is determined by the accuracy of the outcome prediction. $\widehat{\theta}_{\text{PP}}$ will have a small variance when $M\gg m$ to let the second term dominate, and the machine learning predictions are precise enough such that the variance in the second term is small. To see this, when $M\gg m$, we will have $\mathrm{Var}(\widehat{\theta}_{\text{PP}})\approx\frac{1}{m}\mathrm{Var}(Y^{\tt lab}_{i}-\widehat{Y}^{\tt lab}_{i})<\mathrm{Var}(\widehat{\theta}_{\text{CC}})=\frac{1}{m}\mathrm{Var}(Y_{i}^{\tt lab})$ when the predictions are of good quality, or more precisely when the correlation satisfies $\text{Corr}(Y,\widehat{Y})>\frac{1}{2}\sqrt{\mathrm{Var}(Y)/\mathrm{Var}(\widehat{Y})}$.
However, it is critical to understand when $\widehat{{\bm{\theta}}}_{\text{PP}}$ is more efficient than the complete-case analysis in settings other than the mean estimation problem. gronsbell2024another analyzed the efficiency of the $\widehat{{\bm{\theta}}}_{\text{PP}}$ by deriving its influence function under the linear regression setting, and connecting it to the class of AIPW estimators. Specifically, gronsbell2024another shows that $\widehat{{\bm{\theta}}}_{\text{PP}}$ is an AIPW estimator and is not the most efficient in its class. Although the optimal estimator is difficult to obtain, a weighted version of the PPI estimator was proposed to guarantee the efficiency improvement compared to the complete-case analysis (gronsbell2024another). The estimator is of the form $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}_{\text{Chen}}\widehat{\mathbf{\Delta}}$, where $\widehat{\mathbf{W}}_{\text{Chen}}$ is a tuning parameter estimated from the data, originally proposed in chen2000unified under a different double-sampling application context but the same setup. Furthermore, recent modifications to the $\widehat{{\bm{\theta}}}_{\text{PP}}$ all consider multiplying $\widehat{\mathbf{\Delta}}$ with a data-driven weight to seek efficiency improvement over $\widehat{{\bm{\theta}}}_{\text{CC}}$ (angelopoulos2023PPI++; miao2023assumption; gan2024prediction).
@@ 2.4 Limitations of existing prediction-based inference methods
We now discuss the limitations of the existing methods, which we will address. First, existing prediction-based inference methods do not account for the missing-at-random mechanism. For example, in mean estimation, suppose $Y$ is the annual income response in a questionnaire. An auxiliary variable, the level of education, affects the missing probability of $Y$, such that respondents with a lower level of education tend to skip the response to the income question, leaving their $Y$ missing. In such a case, the CCA estimator $\widehat{\theta}_{\text{CC}}$ itself is biased, because it can only reflect the income levels of the population with higher education levels. Second, existing methods only consider missing outcome or missing-multiple-covariate scenarios, although in practice, general missingness patterns commonly appear. For example, in responding to a questionnaire, respondents may randomly skip multiple questions, leading to non-monotone missingness in the resulting dataset; in a longitudinal study, participants may skip a visit and return in the next visit (sun2018inverse). In such cases, existing prediction-based inference methods can only be carried out by removing the incomplete covariates, or truncating a longitudinal trajectory up to when the first missingness occurred.
@@ 3.1 Method
Figure 1: Proposed statistical procedure for ensuring valid prediction-based inference in datasets with general missingness patterns under the missing at random (MAR) assumption. (a) We observe a set of samples with missing values. We assign each missingness pattern with an index; (b) For the $k$-th missingness pattern, we collect the samples with ${\mathcal{C}}_{i}=k$, and impute the missing values with ML predictions. Weighted complete-case analysis is then conducted on the imputed dataset to obtain $\widehat{{\bm{\gamma}}}_{2,k}$; (c) For the completely observed samples, we obtain the weighted analysis estimates $\widehat{{\bm{\theta}}}_{\text{WCC}}$. For each $k$-th missingness pattern, we mask the observed values from the fully observed samples to construct each missingness pattern. We then impute the masked values with machine learning predictions and conduct weighted complete-case analysis to obtain $\widehat{{\bm{\gamma}}}_{1,k}$. $\hat{\Delta}_{k}=\hat{\gamma}_{1,k}-\hat{\gamma}_{2,k}$ converges to zero in probability (implication of Proposition 1).
The PS-PPI estimator is defined as the difference between the WCCA estimator $\widehat{{\bm{\theta}}}_{\text{WCC}}$, computed from the fully observed subset, and a weighted summation of $K$ estimators of zero $\widehat{\mathbf{\Delta}}_{k}$ derived from each missingness pattern:
@@
We introduce each term in (2) in turn. First, $\widehat{{\bm{\theta}}}_{\text{WCC}}$ is the WCCA estimator obtained from the following weighted estimating equations:
$\sum_{i=1}^{N}{\bm{\psi}}^{(i)}_{\text{WCC}}=\sum_{i=1}^{N}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}{\bm{\psi}}\left(\bm{Y}_{i};{\bm{\theta}}\right)=\bm{0},$ (3)
@@
where $\bm{m}$} is a vector-valued function that has the same dimension as $\bm{\omega}$} and satisfies regularity conditions detailed in Section S1.2 of the Supplementary Materials.
Second, $\widehat{\mathbf{\Delta}}_{k}$ is an estimate of zero obtained from the $k$-th missingness pattern. Specifically, we define $\widehat{\mathbf{\Delta}}_{k}=\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k}$, where $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$ are the solutions to the weighted estimating equations:
@@
respectively. That is, we first compute $\widehat{{\bm{\gamma}}}_{1,k}$ by substituting the values in the missing entries with predictions generated from $h_{k}$, and solve the weighted estimating equations based on complete-case data. Second, we compute $\widehat{{\bm{\gamma}}}_{2,k}$ by imputing the actual missing entries with the predictions from $h_{k}$ using data with the $k$-th missingness pattern. We will show that this construction leads to a consistent and asymptotically normal estimator of zero; as shown in Proposition 1, the population parameter of the estimating equations ${\bm{\psi}}_{1,k}$ and ${\bm{\psi}}_{2,k}$ are both ${\bm{\gamma}}^{\ast}_{k}$. We need the following assumption to establish the proposition.
@@
The proof is detailed in Section S1.1 of the Supplementary Materials. The intuition behind Proposition 1 is that, after accounting for the missing data mechanism, the “pseudo” imputed dataset derived from the complete-case data in Equation (5) is distributed as $\mathbb{P}^{(k)}_{\bm{Y}_{\text{imp}}}$, which is identical to the distribution of the imputed dataset based on the subjects with the $k$-th missingness pattern in Equation (6).
Finally, $\{\widehat{\mathbf{W}}_{k}\}$ is a set of tuning parameters in the form of weight matrices with dimension $d\times d$. $\{\widehat{\mathbf{W}}_{k}\}$ may depend on the observed data. In Corollary 3.2, we provide a specific construction that guarantees efficiency gains over WCCA.
@@
We show that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is asymptotically normal.
@@
The proof is detailed in Section S1.2 in the Supplementary Materials. The asymptotic variance-covariance matrix accounts for the uncertainty of the estimated propensity score models. Following sun2018inverse and robins1994estimation, we provide a convenient robust sandwich estimator that ignores the uncertainty contributed by the estimated propensity score model. Such an estimator results in a slightly more conservative yet valid and implementable sandwich variance for $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ when the propensity scores are estimated; the form is shown in Equation (7). Our numerical experiments in Section 4 validate the efficiency advantage of the PS-PPI estimator based on this strategy.
@@
From Equation (7), $\Sigma_{\text{PS-PPI}}$ is a function of $\{\mathbf{W}_{k}\}=(\mathbf{W}_{1},...,\mathbf{W}_{K})$, based on which we may seek the optimal $\{\mathbf{W}^{\ast}_{k}\}$ such that $\Sigma_{\text{PS-PPI}}(\{\mathbf{W}^{\ast}_{k}\})\preceq\Sigma_{\text{PS-PPI}}(\{\mathbf{W}_{k}\})$ for any $\{\mathbf{W}_{k}\}$, where, for two square matrices $\mathbf{A}$ and $\mathbf{B}$, $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite; $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite. Such an optimal set of $\{\mathbf{W}_{k}\}$ has no closed-form and may be computationally expensive to solve. However, when assuming identical $\mathbf{W}_{k}$ across $k$, we have a closed-form solution of $\{\mathbf{W}^{\ast}_{k}\}$, and the resulting $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is guaranteed to be more efficient than $\widehat{{\bm{\theta}}}_{\text{WCC}}$, as we state in Corollary 3.2 and prove in Section S1.3 of the Supplementary Materials.